OpenStudy (anonymous):

hi who discovered e=mc2

OpenStudy (hanah):

first it was discovered by the French physicist/philosopher Poincare, the first man stated the relativity principle, in his equation P=E/C , since P=mv , then at he speed of light mC= E/C, thus E=mc2 . In special relativity the mass energy principle states that in order to preserve the conservation of energy principle we need to presume that mass has an equivalent amount of energy in it represented in the equation m=E/C2 later the relitavistic mechanics , developed by Mincowisci, Planck , Enstien , and many other scientists, proved that there is no physical limit between mass and energy. Read more: http://wiki.answers.com/Q/E_mc2_was_discovered_bY#ixzz23mP74EQF